# Implementing Evidence Acquisition:

### v2

@@ 2 Model
Our model considers an agent (who, depending on the application, may be an individual expert or a team working as a single entity) who can acquire information about an uncertain state $\theta\in\Theta=\{0,1\}$ (e.g., whether the adversary is capable of an attack or whether an aircraft type has a design flaw) at discrete times $\{0,\Delta,2\Delta,\dots,T\}$. For conceptual simplicity, we take $T<\infty$, although our results apply equally to the high-frequency limit as $\Delta\rightarrow 0$. A mechanism designer shares a common prior with the agent over $\theta$; we let $D$ denote the initial probability that $\theta=1$. We have in mind situations where the designer must make some decision at time $T+\Delta$ (e.g., whether to attack the adversary or ground the aircraft type), although as this plays no role we remain agnostic about the designer’s precise preferences—aside from preferring (Blackwell) more information. We first describe the information acquisition technology and then describe the contracting environment.
@@ 2.1 Information Acquisition
The agent can acquire information at time $t$ by paying a cost $c\Delta$, where $c>0$ is an effort cost parameter. When this cost is paid, with some probability a piece of (falsifiable) evidence informative of $\theta$ arrives. We take the probability of evidence arrival to be $\lambda_{\theta}\Delta$; note that this probability may depend on the state. If no evidence arrives, the agent observes a null signal, which we denote by $N$. If the agent does not exert effort, then a null signal is observed with probability 1. Without loss of generality, we assume that $\lambda_{1}\geq\lambda_{0},$ so that the agent’s belief drifts toward state 0 in the absence of evidence arrival. Once the evidence arrives, no further information can be acquired (e.g., only one mechanical error responsible for an aircraft malfunction can be found; if an adversary’s capabilities are determined, no further information is relevant for assessing attack probability, etc.).
When the evidence arrives, the agent observes some $s\in S$—so that, throughout the paper, $S$ is the set of non-null signals. Our main results on static contracts implementing maximal effort relate to the following special cases of this model:
@@
To simplify exposition without detracting from our main message, it suffices to consider the case where $\left\lvert{S}\right\rvert\leq 2$. We refer to non-null signals as either “good news” (e.g., a pilot miscalculation is determined) or “bad news” (e.g., a mechanical issue is found) where a “good news” signal $G$ arrives with probability $\lambda_{\theta}^{G}\Delta$ and a “bad news” signal $B$ arrives with probability $\lambda_{\theta}^{B}\Delta$ when the state is $\theta\in\{0,1\}$ (so that $\lambda_{\theta}=\lambda_{\theta}^{G}+\lambda_{\theta}^{B}$). We take $\lambda_{1}^{G}>\lambda_{0}^{G}$ and $\lambda_{1}^{B}<\lambda_{0}^{B}$. We use the terminology of “bad news” and “good news” to distinguish signals from states, although we do not necessarily view one state as intrinsically better than another (although this may be the case in some settings). Rather, the meaningful difference is that if beliefs drift absent signal arrivals, this drift is toward state 0. As only one piece of evidence can be found, no effort is exerted following the arrival of $G$ or $B$.
This information acquisition technology generalizes Poisson bandit learning (applied to contracting problems by Gerardi and Maestri, (2012); Halac et al., (2016)), as signals (a) need not reveal the state and (b) can be either good or bad. For any $t\leq T$ we let $\mu^{N}_{t}$ denote the posterior belief that $\theta=1$ if no Poisson signal arrived before time $t$ (including $t$) and the agent has exerted effort for all periods until $t$. We refer to $\mu^{N}_{t}$ as the “no-information belief” of the agent that at time $t$. Using the convention that $\mu_{0}^{N}=D$, Bayes rule implies:
@@
noting that signal arrival is off-path whenever $\lambda_{1}^{s}=\lambda_{0}^{s}=0$ (so that beliefs following this event will play no role). Note that if $\lambda_{1}^{G},\lambda_{0}^{G}\in(0,1)$, good news does not reveal the state (similarly for bad news if $\lambda_{1}^{B},\lambda_{0}^{B}\in(0,1)$).
@@ 2.2 Contracting and Main Question
At any time $t$, let $M_{t}$ be the message space of the agent. The history at time $t$ is denoted as $h_{t}=\left\{m_{t^{\prime}}\right\}_{t^{\prime}\leq t}$. Let $\mathcal{H}_{t}$ be the set of all possible histories at time $t$. As our interest is in cases where rewards can condition on the true state ex-post, to avoid trivialization we take rewards to belong to a bounded set, which we take to be the unit interval for simplicity. Therefore, the agent receives a benefit according to the function:
@@
where $R(h_{T},\theta)$ is the fraction of the total available reward provided to the agent when his history of reports is $h_{T}$, and the realized state is $\theta$. An alternative interpretation is that the designer has an indivisible reward (e.g., a special recognition), with $R(h_{T},\theta)$ denoting the probability that the agent receives the reward. [^9] While we allow randomization over the event that the agent receives the full reward, the contracts we defined above are essentially deterministic since the probability that the agent gets the full reward is a deterministic function of the history of reports. In this paper, we focus on deterministic contracts, deferring our discussion of how the possibility of randomization influences the results in Section 6.1. Thus, if the agent has exerted effort in $\tilde{t}$ periods, his final payoff is $R(h_{T},\theta)-c\tilde{t}\Delta$.
Our main findings are qualitatively unchanged if, instead of $\theta$ itself, the designer can condition rewards on some binary signal $\tilde{s}\sim\mathcal{I}(\theta)$ which the agent has no (other) information about. For ease of exposition we do not add this possibility to our model, but in several of our applications such signals may be available even if the state itself is not—for instance, a concrete case is when $\tilde{s}$ is the outcome of some external investigation (e.g., whether the aircraft manufacturer identifies a defect or not).
@@
Scoring rules are substantially simpler compared to arbitrary dynamic contracts, since they only require a single report rather than richer sequences of reports and time-varying rewards. The main goal of our paper is to guide when dynamic reporting is truly necessary to incentivize maximal information acquisition. Simply put, this corresponds to the following question:
@@
We answer this main question using characterizations of effort-maximizing contracts, which also reveal other properties beyond the use of dynamics. These other characteristics may themselves shed light on which incentive schemes best incentivize consultants or forecasters.
@@ 2.3.1 Stopping Strategies
More precisely, call a stopping strategy an information acquisition strategy whereby the agent chooses to (i) exert effort at every time $t\leq\tau$, conditional on the evidence not having yet arrived, and (ii) stop exerting effort once either the signal arrives or $\tau$ has passed. Abusing notation slightly, let $\tau$ denote a stopping time under a stopping strategy.
@@ 2.3.2 Menu Representation
Our next simplification essentially amounts to a version of the taxation principle for our environment. As a stopping strategy involves the agent exerting effort every period until evidence arrives, it is correspondingly not necessary for the agent to report every period. It turns out to be without loss to focus on implementation via a sequence of menu options.
Specifically, we associate each complete sequence of signal realizations—which we think of as the agent’s (endogenously determined) type—with a reward function that maps states into rewards provided to the agent. Note that any type of this form is determined by the length of time the agent exerted effort and the signal observed immediately prior to stopping. We represent a general reward function as $r=(r_{0},r_{1})$, where $r_{0}$ is the reward in state 00 and $r_{1}$ is the reward in state $1$. Our menu representation asserts that it suffices to consider contracts that associate each terminal time-signal pair with a reward function. Thus, each menu option is one such reward function, while the menu offered at time $t$, $\mathcal{R}_{t}$, is the set of reward functions available at time $t$ and after. The agent then makes a one-time, irrevocable choice from the set of available menus.
@@
$u(\mu_{t}^{s},r_{t}^{s})\geq u(\mu_{t}^{s},r),\quad\forall r\in\mathcal{R}_{t},$ (IC)
@@
Section A.1 contains the proof of Lemma 2. Note that as per Eq. IC, given access to the set of contracts in $\mathcal{R}_{t}$, it suffices to consider the incentive constraints where the agent accepts their intended reward function immediately and does not delay. Lemma 2 dramatically simplifies the space of contracts; for instance, it converts the design of the effort-maximizing contracts into a sequence of linear programs for which the sequence of menu options in the effort-maximizing contract can be computed efficiently (see Appendix OA 2 for details).
The above descriptions of incentive compatibility in our menu representation ignore the possibility that the agent stops working prior to either a Poisson signal arrival or time $\tau_{R}$. Of course, this need not be the case. For any $t<\tau_{R}$, we denote
@@
3 Preliminary Intuition and Illustrations
3.1 Preview for Single-Signal Perfect Learning
We preview our findings by presenting the solution in the $\Delta\rightarrow 0$ limit, where a single signal reveals the state, i.e.,
$\lambda^{B}_{0}=\lambda^{B}_{1}=\lambda^{G}_{0}=0,~{}~{}\lambda^{G}_{1}>0.$
We take the horizon $T$ to be sufficiently large so that it will not be a binding constraint. Illustrating this solution will highlight the tensions in maximizing the incentives to exert effort at different points in time.
For this learning environment, we show (in Theorem 2 below) that a scoring rule with two reward functions implements maximum effort: $(1,0)$, corresponding to a guess of state 0, and $(0,r_{1})$, corresponding to a guess of state 1. The value of $r_{1}$, the reward when guessing state 1 (correctly), depends on the initial prior, $D$ (which coincides with $\mu_{0}^{N}$).
Of course, when $\mu_{0}^{N}\in(\mu^{*},\mu^{**})$, if the agent works beliefs may eventually leave this region. If the agent had started at such a belief, more effort could be induced by setting $r_{1}=1$. On the other hand, it is clear why this reoptimization cannot help. When $\mu_{0}^{N}\in(\mu^{*},\mu^{**})$, $r_{1}<1$ will be set so that the initial moral hazard constraint binds—but if the agent expected $r_{1}$ to increase later, he would simply shirk and claim to have found evidence once the reward is increased to $1$. Given that such adjustments are impossible, effort-maximizing mechanisms cannot utilize dynamics and are therefore static.
The lesson is that the tensions between optimizing incentives to exert effort both earlier and later may be unavoidable. A designer may be unable to re-optimize rewards because the re-optimized rewards would violate an incentive constraint at some other time. One deceptive aspect of this example is that the agent’s moral hazard constraint only ever binds at the stopping belief and (possibly) at their initial belief. A technical challenge we face in Section 4.3, for instance, is that if signals are not fully revealing, it may be that the moral hazard constraint binds somewhere “in between.” This feature will imply extra reward functions should be provided to the agent in the effort-maximizing scoring rule. Still, this example illustrates the intuition on how effort-maximizing rewards vary with the agent’s beliefs, which will be useful for understanding the form of effort-maximizing contracts.

### v3

@@ 2 Model
Our model considers an agent (who, depending on the application, may be an individual expert or a team working as a single entity) as a forecaster. The agent is able to acquire information about an uncertain event $\theta\in\Theta=\{0,1\}$ (e.g., whether an investment will achieve some target outcome, whether an adversary will attack on a certain date, etc.) at discrete times $\{\Delta,2\Delta,\dots,T\}$. We refer to $\theta$ as the state. This terminology is consistent with other work on endogenous information acquisition, but we wish to emphasize that we interpret this as a contractible event to be realized in the future (e.g., a successful prototype, an attempted attack from the adversary), following the literature on scoring rule design. For conceptual simplicity, we take $T<\infty$, although our results apply equally to the high-frequency limit as $\Delta\rightarrow 0$. A mechanism designer shares a common prior with the agent over $\theta$; we let $D$ denote the initial probability that $\theta=1$. We have in mind situations where the designer must make some decision at time $T+\Delta$ (e.g., whether to invest in the company or attack the adversary), although as this plays no role we remain agnostic about designer preferences—aside from preferring (Blackwell) more information. We first describe the information acquisition technology and then describe the contracting environment.
@@ 2.1 Information Acquisition
The agent can acquire information at time $t$ by paying a cost $c\Delta$, where $c>0$ is an effort cost parameter. When this cost is paid, with some probability, a piece of evidence arrives that provides information about $\theta$. This evidence is privately observed, and can also be arbitrarily fabricated or misrepresented by the agent at no cost. We refer to this evidence as a Poisson signal. We take the probability of Poisson signal arrival to be $\lambda_{\theta}\Delta$. If no Poisson signal arrives, the agent observes a null signal, which we denote by $N$. If the agent does not exert effort, then a null signal is observed with probability 1. Without loss of generality, we assume that $\lambda_{1}\geq\lambda_{0},$ so that the agent’s belief drifts toward state 0 as no Poisson signal arrives. Once the Poisson signal does arrive, no further information can be acquired (e.g., if a red flag reveals the main point of concern so that additional scrutiny will not meaningfully alter the assessment of the investment’s viability; if an adversary’s capabilities are determined, no further information is relevant for assessing attack probability, etc.). [^7] The assumption of a single-signal arrival represents an extreme case of information attrition in Strulovici, (2022), which discusses various compelling practical instances where the number of available signals is naturally limited. In our case, this feature enhances tractability by allowing us to associate the amount of information produced with the length of time the agent works absent the arrival of a Poisson signal arrival. It also avoids known complications in determining the agent’s payoffs under Poisson learning when signals are not perfectly revealing.
When the Poisson signal arrives, the agent observes some $s\in S$—so that, throughout the paper, $S$ is the set of non-null signals. Our main results on static contracts implementing maximal effort relate to the following special cases of this model:
@@
To simplify exposition without detracting from our main message, it suffices to consider the case where $\left\lvert{S}\right\rvert\leq 2$; doing so provides no restrictions in the second and third environments, and while it does restrict the first environment it turns out our proof for this environment will immediately apply to the case of general $S$ as well. Correspondingly, we refer to non-null signals as either “good news” or “bad news,” where a “good news” signal $G$ arrives with probability $\lambda_{\theta}^{G}\Delta$ and a “bad news” signal $B$ arrives with probability $\lambda_{\theta}^{B}\Delta$ when the state is $\theta\in\{0,1\}$ (so that $\lambda_{\theta}=\lambda_{\theta}^{G}+\lambda_{\theta}^{B}$).
We take $\lambda_{1}^{G}>\lambda_{0}^{G}$ and $\lambda_{1}^{B}\leq\lambda_{0}^{B}$. We use the terminology of “bad news” and “good news” to distinguish signals from states, although we do not necessarily view one state as intrinsically better than another (although this may be the case in some settings). The difference is instead that if beliefs drift absent signal arrivals, this drift is toward state 0. Thus, a “good news” signal could correspond to an undisclosed instance of fraud (i.e., a red flag), while a “bad news” signal might be verification that a key patent is valid (i.e., a green flag). No effort is exerted following the arrival of $G$ or $B$, as doing so reveals no further information.
This information acquisition technology generalizes Poisson bandit learning (applied to contracting problems by Gerardi and Maestri, (2012); Halac et al., (2016)), as signals (a) need not reveal the state and (b) can be either good or bad. For any $t\leq T$ we let $\mu^{N}_{t}$ denote the posterior belief that $\theta=1$ if no Poisson signal arrived before time $t$ (including $t$) and the agent has exerted effort for all periods until $t$. We refer to $\mu^{N}_{t}$ as the “no-information belief” of the agent at time $t$. Using the convention that $\mu_{0}^{N}=D$, Bayes rule implies:
@@
noting that signal arrival is off-path whenever $\lambda_{1}^{s}=\lambda_{0}^{s}=0$. Note that if $\lambda_{1}^{G},\lambda_{0}^{G}\in(0,1)$, good news does not reveal the state (similarly for bad news if $\lambda_{1}^{B},\lambda_{0}^{B}\in(0,1)$).
@@ 2.2 Contracting and Main Question
Let $M_{t}$ be the message space of the agent at any time $t$. We denote the history at time $t$ as $h_{t}=\left\{m_{t^{\prime}}\right\}_{t^{\prime}\leq t}$. Let $\mathcal{H}_{t}$ be the set of all possible histories at time $t$. Since we allow rewards to condition on the future realized outcome of interest (i.e., the state), we take rewards to belong to a bounded set, which we take to be the unit interval for simplicity. As mentioned in the introduction, we interpret the reward as an endorsement which positively influences the forecaster’s reputation. Such reward bounds are also widely assumed in the literature on evaluating forecasters (e.g., Deb et al., , 2018), as well as the literature on information elicitation more generally. [^8] While we interpret the reward as non-monetary, our model also applies to applications where the designer faces an explicit budget constraints set by a third party. Anthony et al., (2007) discusses firms imposing budget constraints on different divisions; National meteorological agencies, such as the U.S. National Weather Service (NWS) receive public funding for operations; the Congressional Budget Office (CBO) operates under a fixed annual budget and provides forecasts for Congress. In such cases, designers are restricted to using this budget to incentivize information acquisition. The agent thus receives a benefit according to the function:
@@
where $R(h_{T},\theta)$ is the fraction of the total available reward provided to the agent when his history of reports is $h_{T}$, and the realized state is $\theta$—or, alternatively, the probability that the agent receives the reward, so that $R(h_{T},\theta)$ can be interpreted as an endorsement probability. [^9] While we allow randomization over the event that the agent receives the full reward, we do not allow stochastic messages from the principal to the agent. A discussion of how the possibility of such randomization influences the results is deferred to Section 6.1. Allowing stochastic messages would introduce a second wedge between scoring rules and arbitrary contracts, which we do not consider to maintain focus on the ability to elicit information over time. If the agent has exerted effort in $\tilde{t}$ periods, his final payoff is $R(h_{T},\theta)-c\tilde{t}\Delta$.
@@
Scoring rules are substantially simpler compared to arbitrary dynamic contracts, since they only require a single report rather than richer sequences of reports and time-varying rewards. As discussed in the introduction, contracts that facilitate richer reporting capabilities may require the expert to be more closely integrated within the decision maker’s organization. The main goal of our paper is to characterize whether unrestricted dynamic reporting is truly necessary to incentivize maximal information acquisition, or whether a report summarizing the acquired information suffices. Simply put, this corresponds to the following question:
@@
We answer this main question using characterizations of effort-maximizing contracts, which also reveal other properties beyond the use of dynamics. These other characteristics may be of independent interest as they shed light on which incentive schemes best incentivize consultants or forecasters.
@@ 2.3.1 Stopping Strategies
More precisely, call a stopping strategy an information acquisition strategy whereby the agent chooses to (i) exert effort at every time $t\leq\tau$, conditional on no Poisson signal having yet arrived, and (ii) stop exerting effort once either the Poisson signal arrives or $\tau$ has passed. Abusing notation slightly, let $\tau$ denote a stopping time under a stopping strategy.
@@ 2.3.2 Menu Representation
Our next simplification essentially amounts to a version of the taxation principle for our environment. As a stopping strategy involves the agent exerting effort every period until a Poisson signal arrives, it turns out to be without loss to focus on implementation via a sequence of menu options.
Specifically, we associate each complete sequence of signal realizations—which we think of as the agent’s (endogenously determined) type—with a reward function that maps states into rewards (e.g., endorsement probabilities) provided to the agent. Note that any type of this form is determined by the length of time the agent exerted effort and the signal observed immediately prior to stopping. We represent a general reward function as $r=(r_{0},r_{1})$, where $r_{0}$ is the reward in state 00 and $r_{1}$ is the reward in state $1$. Our menu representation asserts that it suffices to consider contracts that associate each terminal time-signal pair with a reward function. Thus, each menu option is one such reward function, while the menu offered at time $t$, $\mathcal{R}_{t}$, is the set of reward functions available at time $t$ and after. The agent then makes a one-time, irrevocable choice from the set of available menus.
@@
$u(\mu_{t}^{s},r_{t}^{s})\geq u(\mu_{t}^{s},r),\quad\forall t\leq T,s\in S,r\in\mathcal{R}_{t},$ (IC)
@@
Appendix A contains the proof of Lemma 2. Note that as per Eq. IC, given access to the set of contracts in $\mathcal{R}_{t}$, it suffices to consider the incentive constraints where the agent accepts their intended reward function immediately and does not delay. Lemma 2 dramatically simplifies the space of contracts; it converts the design of the effort-maximizing contracts into a sequence of linear programs whose solution can be computed efficiently, as we illustrate in Section 3.1.
The above descriptions of truthful reporting in our menu representation ignore the possibility that the agent stops working prior to either a Poisson signal arrival or time $\tau_{R}$. Of course, this need not be the case. For any $t<\tau_{R}$, we denote
@@
3 Preliminary Observations
3.1 Numerical Illustration of Effort-Maximizing Contracts
We describe a sequence of linear programs which can be used to identify an effort-maximizing contract. One simple observation is that as long as the environment is non-stationary, it is impossible to implement effort in every period if $c>0$. Indeed, for non-stationary environments $\mu^{N}_{t}\rightarrow 0$ as $t\rightarrow\infty$. And if $\mu^{N}_{t}<c$, it is not possible to incentivize any effort—even if the agent’s payoff increased by 1 in state 1 when exerting effort, this would not suffice to cover its cost. While it is not necessarily the case that effort can be incentivized whenever $\mu^{N}_{t}>c$, as $c\rightarrow 0$, the maximum time the agent would work under a contract that pays a reward if and only if the agent guesses the state correctly approaches infinity in this limit. [^11] For environments where good news reveals the state, this can be achieved via a contract that pays a reward of 1 when the state is guessed correctly. If not, then the reward when guessing state 0 must decrease as both $c$ and the stopping belief approaches 0, in order to ensure that different guesses are made following different signal realizations.
Given these observations, fixing the learning technology and a horizon $\tilde{t}$, there exists some cost $c^{*}(\tilde{t})$ such that a contract $R$ exists with $\tau_{R}\geq\tilde{t}$ whenever $c\leq c^{*}(\tilde{t})$ but not whenever $c>c^{*}(\tilde{t})$. In order to compute an effort-maximizing contract given cost $c$, we first solve the problem of verifying whether there is an effort-maximizing contract implementing effort until some time $\tilde{t}$ given $c$. We do this by computing the maximum cost $c^{*}(\tilde{t})$ such that the agent can be incentivized to exert effort until time $\tilde{t}$, and then checking whether $c\leq c^{*}(\tilde{t})$. Then, an effort maximizing contract is one such that, for some $t^{*}\geq 0$, $c^{*}(t^{*})\geq c>c^{*}(t^{*}+1)$ (and if no such $t^{*}$ exists, effort cannot be implemented).
Given any $\tilde{t}$, the maximum effort $c=c^{*}(\tilde{t})$ is the solution of the following program (where we recall that $\mathcal{R}_{t}^{\tilde{t}}=\{r^{s}_{t^{\prime}}\}_{t\leq t^{\prime}\leq\tilde{t},s\in S}\cup\{r^{N}_{\tilde{t}}\}$):
$\max_{\tilde{c},\mathcal{R}_{0}^{\tilde{t}}}\hskip 20.0pt$ $~{}~{}~{}~{}~{}~{}\tilde{c}$
s.t. $V_{t}(\mu_{t}^{N})\geq u(\mu_{t}^{N},\tilde{r}),$ $\forall t\in\{0,\ldots,\tilde{t}\},\tilde{r}\in\mathcal{R}_{t}^{\tilde{t}}$
$u(\mu_{t}^{s},r_{t}^{s})\geq u(\mu_{t}^{s},\tilde{r}),$ $\forall t\in\{0,\ldots,\tilde{t}\},\tilde{r}\in\mathcal{R}_{t}^{\tilde{t}},s\in\{G,B\}$
$u(\mu_{\tilde{t}}^{N},r_{\tilde{t}}^{N})\geq u(\mu_{\tilde{t}}^{N},r_{\tilde{t}}^{s}),$ $\forall s\in\{G,B\}$
$0\leq r_{t}^{s}(\theta)\leq 1,$ $\forall t\in\{0,\ldots,\tilde{t}\},s\in\{G,B\}$
$0\leq r_{\tilde{t}}^{N}(\theta)\leq 1$
where $V_{t}(\mu_{t}^{N})$ is the agent’s expected payoff under the contract when exerting effort until time $\tilde{t}$ in the absence of a Poisson signal; i.e., letting $f^{s}_{t}(t^{\prime})$ be the probability of receiving Poisson signal $s$ at time $t^{\prime}$ conditional on not receiving Poisson signals before time $t$ and $F^{s}_{t}(t^{\prime})$ be the corresponding cumulative probability:
$V_{t}(\mu_{t}^{N})=\left(1-\sum_{s\in S}F^{s}_{t}(\tilde{t})\right)\cdot(u(\mu^{N}_{\tilde{t}},r^{N}_{\tilde{t}})-c\cdot(\tilde{t}-t))+\sum_{t^{\prime}=t}^{\tilde{t}}\sum_{s\in S}(u(\mu^{s}_{t^{\prime}},r^{s}_{t^{\prime}})-c\cdot(t^{\prime}-t))\cdot f^{s}_{t}(t^{\prime}).$
In particular, by inspection we see that $\tilde{c}$ as well each reward function enters the constraints linearly, since $u(\mu,r)$ is linear in $r$. Using standard linear programming algorithms, this formulation delivers a way of numerically calculating effort-maximizing contracts.
Time
Time
Figure 1: Solution to the linear program for the effort minimizing contract/maximal possible effort implementing effort for 10 periods, allowing for reporting at any time. Parameters chosen are $\lambda_{1}^{G}=1/3,\lambda_{1}^{B}=\lambda_{0}^{G}=1/10,\lambda_{0}^{B}=1/5,D=2/3$.
Figure 1 presents the solution to this linear program for certain parameter values. In Section 5, we show that the declining reward structure exhibited by this solution characterizes effort-maximizing contracts when reporting signal $B$ in general. We also characterize the rewards provided when reporting signal $G$ (Theorem 5). On the other hand, for a given cost level and arbitrary parameters, there will typically not be a unique effort-maximizing contract, as slight adjustments in the rewards could also yield the same level of implemented effort without violating incentive compatibility or influencing effort incentives. Nevertheless, this procedure provides a numerical calculation of an effort-maximizing contract.
Importantly, the above linear program allows the principal to discriminate on time. An agent who observes a signal $s$ at time $t$ cannot choose the reward function that would have been selected had that signal arrived at time $t^{\prime}<t$. However, if the principal were to utilize a single-elicitation mechanism, such deviations would be possible. Therefore, to identify effort-maximizing contracts featuring summarized advice—i.e., a scoring rule—we can instead follow the same produced as we followed to identify effort-maximizing (dynamic) contracts, using the following linear program instead:
$\max_{\tilde{c},\mathcal{R}_{0}^{\tilde{t}}}\hskip 15.0pt$ $~{}~{}~{}~{}\tilde{c}$
s.t $V_{t}(\mu_{t}^{N})\geq u(\mu_{t}^{N},\tilde{r}),$ $\forall t\in\{0,\ldots,\tilde{t}\},\tilde{r}\in\mathcal{R}_{0}^{\tilde{t}}$
$u(\mu_{t}^{s},r_{t}^{s})\geq u(\mu_{t}^{s},\tilde{r}),$ $\forall t\in\{0,\ldots,\tilde{t}\},\tilde{r}\in\mathcal{R}_{0}^{\tilde{t}},s\in\{G,B\}$
$u(\mu_{\tilde{t}}^{N},r_{\tilde{t}}^{N})\geq u(\mu_{\tilde{t}}^{N},\tilde{r}),$ $\forall\tilde{r}\in\mathcal{R}_{0}^{\tilde{t}}$
$0\leq r_{t}^{s}(\theta)\leq 1,$ $\forall t\in\{0,\ldots,\tilde{t}\},s\in\{G,B\}$
$0\leq r_{\tilde{t}}^{N}(\theta)\leq 1$
We illustrate the solution to this linear program in Figure 2. Here, different reward functions at different times correspond to what the agent selects if the signal arrives at that time, even though no report is made until after time 10 (since the agent’s belief will differ depending on how long it takes for the signal to arrive). Indeed, we see that this contract is in fact different than the one involving dynamic elicitation.
While not obvious from the figures alone, it turns out that it is indeed the case for this informational environment that there is a range of costs such that the principal can implement effort to time 10 under a dynamic contract, but not using static elicitation (specifically, whenever $0.0408335\leq c\leq 0.0494949$). Our main results precisely elucidate why this is the case, showing that we would obtain no gap in the single-signal, stationary or perfectly revealing environments—the parameters for these solutions, by contrast, do not fit under any of these environments.
One natural way to try to prove the sufficiency of summarized advice would be to show that the constraints added to the second program do not bind under these assumptions. Unfortunately, we do not see a clear way of determining which constraints will bind at which time given properties of the linear program alone.
To summarize, while the numerical calculation shows how to identify effort-maximizing contracts given an information acquisition technology, to our knowledge they are of limited use for determining whether a static implementation exists of the effort maximizing contract. The illustrations suggest doing so is nontrivial, since we can identify parameters such that there is indeed a gap between the amount of effort implementable across different contracting assumptions. Hence, the exercise motivates us to adopt different techniques to determine the qualitative properties of such contracts, which we view as economically significant.
Time
Time
Figure 2: Solution to the linear program for the effort minimizing contract/maximal possible effort implementing effort for 10 periods, assuming that the contract involves a single elicitation at time 10. Parameters chosen are $\lambda_{1}^{G}=1/3,\lambda_{1}^{B}=\lambda_{0}^{G}=1/10,\lambda_{0}^{B}=1/5,D=2/3$.